OpenStudy (anonymous):

If f(6)=9 and f(x)=g(x)+2, what point can you determine on the graph of g? Guys, how do we solve this?

OpenStudy (dayakar):

f(6) = 9-----(1) f(x) = g(x) +2 g(x) = f(x) -2 -------(2) if x=6 g(6) = f(6) -2 = 9 -2 [ from (1) f(6) =9] do remaining part

OpenStudy (dayakar):

do u understand this

OpenStudy (anonymous):

@dayakar Hey! Thank you. But how did 2 become negative?

OpenStudy (anonymous):

@dayakar oh, because we substracted g(x) from both sides in the equation?

OpenStudy (skullpatrol):



OpenStudy (anonymous):

@skullpatrol Great! Thanks! Yay

OpenStudy (skullpatrol):

Thanks for asking :D

OpenStudy (skullpatrol):

The goal is to get g(x) alone on one side of the equation and then use the given information.

OpenStudy (anonymous):

@skullpatrol that's been such a great help!

OpenStudy (skullpatrol):

That's great to hear :-)